Canada’s real GDP was $1,757.6 billion in 2014 and $1762.2 billion in 2015. Canada’s population growth rate in 2015 was 0.8 percent. Calculate:

2a. Canada’s economic growth rate in 2015.

2b. The growth rate of real GDP per person in Canada in 2015.

2c. The approximate number of years it takes for real GDP per person in Canada to double if the 2015 economic growth rate and population growth rate are maintained.

2d. The approximate number of years it takes for real GDP per person in Canada to double if the economic growth rate rises to 6 percent a year but the population growth rate remains the same as it was in 2015.
Answer #1

2a. Growth Rate = Change in GDP/GDP in 2014

= (1762.2 - 1757.6)/1757.6 = 0.26%

2b. We assume a population base of 100 for 2014.

GDP per person in 2014 = 1757.6/100 = 17.576

  GDP per person in 2015 = 1762.2/100*(1+ 0.8%) = 17.482

Growth Rate = Change in GDP per person /GDP per person in 2014

= (17.482 - 17.576)/17.576 = -0.53%

2c. GDP growth rate = 0.26%

Population growth rate = 0.8%

Since the population growth is higher than the GDP growth rate, GDP per person will decline and never double.

2d. GDP growth rate = 6%.....(i)

Population growth rate = 0.8%....(ii)

Assuming GDP = x and population = y and using (i)& (ii)

Growth Rate = Change in GDP per person /GDP per person in 2015

= (1.06x/1.008y - x/y)/ (x/y)

= 5.16%

Let it take n yerars to double,

FV = PV(1 + r)^n

2 = (1 + 5.16%)^n

Solving, n ~ 14 years